OpenStudy (anonymous):

Television sets. What does it mean to refer to a 20 in TV set? Such units refer to the diagonal of the screen. A 15 in TV set also has a width of 12 inches. What is its height? ___ inches

OpenStudy (anonymous):

its a 12 by x by 15. plug into 12^2+x^2=15^2 and you get x=9

OpenStudy (anonymous):

\[12^{2}+x ^{2}=15^{2}\] x=9